# Animate3D: Animating Any 3D Model with Multi-view Video Diffusion

### v1

@@ Abstract
Recent advances in 4D generation mainly focus on generating 4D content by distilling pre-trained text or single-view image-conditioned models. It is inconvenient for them to take advantage of various off-the-shelf 3D assets with multi-view attributes, and their results suffer from spatiotemporal inconsistency owing to the inherent ambiguity in the supervision signals. In this work, we present Animate3D, a novel framework for animating any static 3D model. The core idea is two-fold: 1) We propose a novel multi-view video diffusion model (MV-VDM) conditioned on multi-view renderings of the static 3D object, which is trained on our presented large-scale multi-view video dataset (MV-Video). 2) Based on MV-VDM, we introduce a framework combining reconstruction and 4D Score Distillation Sampling (4D-SDS) to leverage the multi-view video diffusion priors for animating 3D objects. Specifically, for MV-VDM, we design a new spatiotemporal attention module to enhance spatial and temporal consistency by integrating 3D and video diffusion models. Additionally, we leverage the static 3D model’s multi-view renderings as conditions to preserve its identity. For animating 3D models, an effective two-stage pipeline is proposed: we first reconstruct motions directly from generated multi-view videos, followed by the introduced 4D-SDS to refine both appearance and motion. Qualitative and quantitative experiments demonstrate that Animate3D significantly outperforms previous approaches. Data, code, and models will be open-released.
@@ 1 Introduction
3D content creation has garnered significant attention due to its wide applicability in AR/VR, gaming, and movie industry. With the development of diffusion models [48, 47, 42, 20, 54, 6, 19] and large-scale 3D object datasets [13, 12, 39, 69, 74, 14], recent 3D foundational generations have seen extensive exploration through fine-tuned text-to-image (T2I) diffusion models [33, 43, 44, 55, 34, 52], as well as training large reconstruction models from scratch [22, 64, 50, 72, 57], leading the 3D assets creation to a new era. Despite the significant progress in static 3D representation, this momentum has not been paralleled in the realm of dynamic 3D content generation, also known as 4D generation.
@@
To further enable impressive animations from 3D objects that can be observed at any viewpoint and time, we jointly optimize the 4DGS [58] through both reconstruction and 4D Score Distillation Sampling (4D-SDS) losses based on our unified MV-VDM. Benefiting from the spatiotemporal consistent multi-view video generations, 4DGS can be roughly converged to proper results with only reconstruction loss, while 4D-SDS further improves the details and fine-grained motions.
@@
We highlight the contribution of this paper as follows: 1) Animate3D is the first 4D generation framework to animate any 3D objects with detailed multi-view conditions. 2) We propose the foundational 4D generation model, MV-VDM, to jointly model spatiotemporal consistency. 3) We present the largest high-quality 4D datasets MV-Video collected with about 115K animations with over 1.8M multi-view videos. Extensive experiments demonstrate that our data-driven approach can generate spatiotemporal consistent 4D objects, significantly outperforming previous counterparts.
@@ 4.1 Setup
Implementation Details. We sample 16 frames evenly for each animation to train our MV-VDM. We use the Adamw optimizer with a learning rate of $42e-$ 4 and a weight decay $0.01$, and train the model for 20 epochs with a batch size of 1024. When inference, we set the sampling step to 25 and adopt freeinit [59] to get stable results when animating 3D objects. As for 4D generation, the resolution and feature dimension of the Hex-planes are set to $[100,100,8]$ and 16, respectively. We perform motion reconstruction in progressive way for the first 750 iterations with a batch size of 64 (4 views, 16 frames), and then add 4D-SDS optimization for another 250 iterations. Learning rate is 0.01 and 0.0001 for Hex-plane and offset prediction layers. $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ in Eq. 11 are set to $100.0$, $0.01$ and $10.0$ , respectively. It costs 10 days to train our MV-VDM on 32 80G A800 GPUs, and the optimization for 4D generation takes around 40 minutes (20 minutes for motion reconstruction, and 20 minutes for 4D-SDS) on a single A800 GPU per object.

### v2

@@ Abstract
Recent advances in 4D generation mainly focus on generating 4D content by distilling pre-trained text or single-view image-conditioned models. It is inconvenient for them to take advantage of various off-the-shelf 3D assets with multi-view attributes, and their results suffer from spatiotemporal inconsistency owing to the inherent ambiguity in the supervision signals. In this work, we present Animate3D, a novel framework for animating any static 3D model. The core idea is two-fold: 1) We propose a novel multi-view video diffusion model (MV-VDM) conditioned on multi-view renderings of the static 3D object, which is trained on our presented large-scale multi-view video dataset (MV-Video). 2) Based on MV-VDM, we introduce a framework combining reconstruction and 4D Score Distillation Sampling (4D-SDS) to leverage the multi-view video diffusion priors for animating 3D objects. Specifically, for MV-VDM, we design a new spatiotemporal attention module to enhance spatial and temporal consistency by integrating 3D and video diffusion models. Additionally, we leverage the static 3D model’s multi-view renderings as conditions to preserve its identity. For animating 3D models, an effective two-stage pipeline is proposed: we first reconstruct motions directly from generated multi-view videos, followed by the introduced 4D-SDS to refine both appearance and motion. Benefiting from accurate motion learning, we could achieve straightforward mesh animation. Qualitative and quantitative experiments demonstrate that Animate3D significantly outperforms previous approaches. Data, code, and models will be open-released.
@@ 1 Introduction
3D content creation has garnered significant attention due to its wide applicability in AR/VR, gaming, and movie industry. With the development of diffusion models [48, 47, 42, 20, 54, 6, 19] and large-scale 3D object datasets [13, 12, 39, 69, 74, 14], recent 3D foundational generations have seen extensive exploration through fine-tuned text-to-image (T2I) diffusion models [33, 43, 75, 44, 55, 34, 52], as well as training large reconstruction models from scratch [22, 64, 50, 72, 57], leading the 3D assets creation to a new era. Despite the significant progress in static 3D representation, this momentum has not been paralleled in the realm of dynamic 3D content generation, also known as 4D generation.
@@
To further enable impressive animations from 3D objects that can be observed at any viewpoint and time, we jointly optimize the 4DGS [58] through both reconstruction and 4D Score Distillation Sampling (4D-SDS) losses based on our unified MV-VDM. Benefiting from the spatiotemporal consistent multi-view video generations, 4DGS can be roughly converged to proper results with only reconstruction loss, while 4D-SDS further improves the details and fine-grained motions. The Gaussian trajectory learned by our framework is surprisingly accurate and could be used to directly animate the mesh.
@@
We highlight the contribution of this paper as follows: 1) Animate3D is the first 4D generation framework to animate any 3D objects with detailed multi-view conditions. The framework is further extended to achieve mesh animation without skeleton rigging. 2) We propose the foundational 4D generation model, MV-VDM, to jointly model spatiotemporal consistency. 3) We present the largest high-quality 4D datasets MV-Video collected with about 115K animations with over 1.8M multi-view videos. Extensive experiments demonstrate that our data-driven approach can generate spatiotemporal consistent 4D objects, significantly outperforming previous counterparts.
@@ 3.2 Reconstruction and Distillation of 4DGS
3.3 Extension to Mesh Animation
To directly utilize high-quality mesh generated from commercial 3D generation tools and crafted by human experts, we extent our framework to mesh animation, producing animated mesh compatible with standard 3D rendering pipelines.
We initialize the 3DGS representation of the given object by vertices and triangles of the static mesh. Specifically, the color is determined by vertex color and we average the connected edges for the scale. Opacity and rotation are set to fully visible and zero rotation quaternion, respectively. The coarse 3DGS is animated following the motion reconstruction steps as described in the above sections. We utilize the per-vertex Gaussian trajectory to deform the static mesh in a straightforward way without skeleton rigging, control point selection or complicated deformation algorithms. As shown in Fig. 5 and our project page, the results are surprisingly good despite the simplicity of the solution.
@@ 4.1 Setup
Implementation Details. We sample 16 frames evenly for each animation to train our MV-VDM. We use the Adamw optimizer with a learning rate of $42e-$ 4 and a weight decay $0.01$, and train the model for 20 epochs with a batch size of 1024. When inference, we set the sampling step to 25 and adopt freeinit [59] to get stable results when animating 3D objects. As for 4D generation, the resolution and feature dimension of the Hex-planes are set to $[100,100,8]$ and 16, respectively. We perform motion reconstruction in progressive way for the first 750 iterations with a batch size of 64 (4 views, 16 frames), and then add 4D-SDS optimization for another 250 iterations. Learning rate is 0.01 and 0.0001 for Hex-plane and offset prediction layers. $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ in Eq. 11 are set to $100.0$, $0.01$ and $10.0$ , respectively. It costs 10 days to train our MV-VDM on 32 80G A800 GPUs, and the optimization for 4D generation takes around 40 minutes (20 minutes for motion reconstruction, and 20 minutes for 4D-SDS) on a single A800 GPU per object. Mesh animation takes 15 minutes, faster than standard 3DGS animation because it uses fewer points in 3DGS and only requires motion reconstruction stage.
@@ 4.3 Ablation
4.4 Mesh Animation
Figure 5: Visualizations of mesh animation. We present RGB and textureless renderings of two mesh animation results. Best viewed by zooming in.
We provide mesh animation results in Fig. 5. Static meshes are generated by commercial 3D generation tools. For more results, please visit our project page.